OpenStudy (anonymous):

Pam is looking for a new car. She has test-driven two cars but can only purchase one. The probability that she will purchase car A is 0.46, and the probability that she will purchase car B is 0.34. What is the probability that she will not purchase either car A or car B? 0.12 0.40 0.20 0.80 ________________________________________

OpenStudy (mathstudent55):

The sum of all probabilities is 1.

OpenStudy (igreen):

Right.

OpenStudy (igreen):

0.46 + 0.34 = 0.80 So what is the probability she won't choose any of the cars? (Hint: Subtract from 1)

OpenStudy (mathstudent55):

p(buy car A) + p(buy car B) + p(buy no car) = 1 Replace the two known probabilities by their numbers, and solve for p(buy no car).

OpenStudy (igreen):

Can you do that? @MACH1303

OpenStudy (anonymous):

yes give me a sec

OpenStudy (igreen):

\(1 - 0.8 =~?\)

OpenStudy (anonymous):

.2 my computer was messing with me thnx

OpenStudy (igreen):

No problem. That's correct. :)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. \(\bbox[5pt, #000000 ,border:5px solid #00FF00]{\bf\huge \color {#00FF00}\sim \! \! \! \! \color{#00FF00}i\it\Huge\color{#00FF00}G \color{#00FF00}r\color{#00FF00}e\color{#00FF00}e\color{#00FF00}n\bf \color{#00FF00}ツ}\)